OpenStudy (anonymous):

x = 2t – 1 y = t2 + 5, -4 ≤ t ≤ 4

OpenStudy (anonymous):

i need to graph these

OpenStudy (anonymous):

need to be graphed and im lost

OpenStudy (ankit042):

hmm y=t^2+5 and t =(x+1)/2 replace this in the first equation to continue

OpenStudy (anonymous):

so would it be (x+1)/2 + 5 ?

OpenStudy (ankit042):

No you have to square t

OpenStudy (anonymous):

(x+1)/(2 + 5))^2 ?

OpenStudy (anonymous):

?????????